# Brightest Cluster Galaxy Offsets in Cold Dark Matter

### v1

@@ Abstract
The distribution of offsets between the brightest cluster galaxies of galaxy clusters and the centroid of their dark matter distributions is a promising probe of the underlying dark matter physics. In particular, since this distribution is sensitive to the shape of the potential in galaxy cluster cores, it constitutes a test of dark matter self-interaction on the largest mass scales in the universe. We examine these offsets in three suites of modern cosmological simulations; IllustrisTNG, MillenniumTNG and BAHAMAS. For clusters above $10^{14}\,\rm{M_{\odot}}$, we examine the dependence of the offset distribution on gravitational softening length, the method used to identify centroids, redshift, mass, baryonic physics, and establish the stability of our results with respect to various nuisance parameter choices. We find that offsets are overwhelmingly measured to be smaller than the minimum converged length scale in each simulation, with a median offset of $\sim 1\,\rm{kpc}$ in the highest resolution simulation considered, TNG300-1, which uses a gravitational softening length of $1.48\,\rm{kpc}$. We also find that centroids identified via source extraction on smoothed dark matter and stellar particle data are consistent with the potential minimum, but that observationally relevant methods sensitive to cluster strong gravitational lensing scales, or those using the the “light traces mass” approach, in this context meaning gas is used as a tracer for the potential, can overestimate offsets by factors of $\sim 10$ and $\sim 30$, respectively. This has the potential to reduce tensions with existing offset measurements which have served as evidence for a nonzero dark matter self-interaction cross section.
keywords:
cosmology: dark matter – galaxies: clusters – gravitational lensing
pubyear: 2024pagerange: Brightest Cluster Galaxy Offsets in Cold Dark MatterBrightest Cluster Galaxy Offsets in Cold Dark Matter
@@ 1 Introduction
The first evidence for the existence of a dominant and invisible matter component in the universe came from observations of the Coma galaxy cluster in 1933 (Zwicky, 2009), for which the kinematics of the galaxies within suggested an average density 400 times the inferred density from visible matter. The evidence for this dark matter (DM) has since accumulated to include for example the rotation curves of galaxies (Rubin & Ford, 1970; Ou et al., 2023), their escape velocities (Necib & Lin, 2022), the cosmic microwave background (CMB; Planck Collaboration et al. 2020), and gravitational lens modelling (Clowe et al., 2006; Sharon et al., 2020). The particle nature of DM remains one of the primary questions of modern astrophysics, but a cold, collisionless DM in tandem with a cosmological constant $\Lambda$ constitutes the current standard cosmological model (Weinberg et al., 2013; Vogelsberger et al., 2020). The $\Lambda$ cold dark matter (hereafter $\Lambda$CDM) cosmological paradigm has been remarkably successful in explaining the large scale structure of the universe via the Lyman-alpha forest (Evslin, 2017; Addison et al., 2018; Cuceu et al., 2019), the matter power spectrum and galaxy clustering (Reid et al., 2010; Eisenstein et al., 2011), the accelerated expansion of the universe (Riess et al., 1998; Perlmutter et al., 1999) and observations of the cosmic microwave background anisotropies and polarizations (Page et al., 2003).
@@
Galaxy clusters form at the densest nodes in the cosmic web, with the highest dark matter densities in their cores. These cores constitute ideal testing grounds to search for cored dark matter profiles, for example via strong lens modelling (Zitrin et al., 2012; Limousin et al., 2022), the internal kinematics of the brightest cluster galaxy (BCG; Kelson et al. 2002), weak lensing, and combinations of these probes (Newman et al., 2013). Evidence for both cored and cuspy cluster core profiles has been discovered using these methods (Biviano et al., 2023), with the source of the measured coring not yet clearly attributable to baryonic feedback or dark matter particle interactions alone. This uncertainty is due in part to the fact that baryonic effects have been demonstrated to both increase DM density slopes in the cores of clusters via adiabatic contraction (Schaller et al., 2015a), and decrease these slopes via dynamical friction (Inoue & Saitoh, 2011; Martizzi et al., 2012), galaxy formation processes (Laporte et al., 2012) and active galactic nucleus feedback (Ragone-Figueroa et al., 2012; Peirani et al., 2017).
@@
BAHAMAS 400 $1024^{3}$ $8.1\times 10^{8}$ $4.5\times 10^{9}$ 4 4 2624 30
Table 1: Physical parameters of the simulations considered in this article. Parameters are (from left to right): the simulation identifier, length of one side of the cubic simulation box, $N_{\rm{part}}$ is both the initial number of gas particles (some of which will be converted to stellar particles as time evolves), and also the number of DM particles (which remains unchanged), initial baryonic mass for moving-mesh code $M_{\rm{b}}$, dark matter particle mass $M_{\rm{DM}}$, gravitational softening length of the dark matter and star particles $\epsilon_{\rm{DM,\star}}$ at $z=0$, minimum gravitational softening length of the adaptive gas softening $\epsilon_{\rm{gas,min}}$ in comoving coordinates, number of friends-of-friends identified groups with mass above $10^{14}\,\rm{M_{\odot}}$ at $z=0$, and number of friends-of-friends identified groups with mass above $10^{15}\,\rm{M_{\odot}}$ at $z=0$.
@@ 2.1 IllustrisTNG
The IllustrisTNG project (TNG; Springel et al. 2017; Pillepich et al. 2017b; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2021) is a suite of cosmological hydrodynamical simulations spanning various simulation box sizes, mass resolutions, and physics models based on the AREPO code (Springel, 2010). This project succeeds the Illustris project (Vogelsberger et al., 2014a; Genel et al., 2014; Vogelsberger et al., 2014b; Sijacki et al., 2015), improving upon the galaxy formation model (Vogelsberger et al., 2013) and expanding the scope to larger volumes and higher resolutions. The TNG simulation suite consists of TNG50, TNG100 and TNG300, characterized by their varying simulation box sizes of $35$, $75$ and $205\,\rm{cMpc}$$\,h^{-1}$ respectively.
@@ 2.2 MillenniumTNG
The MillenniumTNG project (MTNG; Hernández-Aguayo et al. 2023; Pakmor et al. 2023; Barrera et al. 2023) is the successor to the dark matter-only Millennium simulation (Springel et al., 2005), the first 10 billion particle simulation in a $500\,\rm{Mpc}$$\,h^{-1}$ box. MillenniumTNG combines the volume of Millennium with the hydrodynamical modelling [^2] With some small fixes in addition to changes relative to the TNG model, implemented to save memory in the much larger box (Pakmor et al., 2023). developed for the IllustrisTNG project (Weinberger et al., 2016; Pillepich et al., 2017a) and is run using an improved version of the AREPO code (Pakmor et al., 2015; Weinberger et al., 2020). The MillenniumTNG simulations involve high time resolution merger trees and direct light cone output, enabling detailed cosmological inference via comparisons to large scale surveys (Contreras et al., 2023; Kannan et al., 2023; Bose et al., 2023) and studies of the impact of baryons and massive neutrinos on weak lensing observations (Ferlito et al., 2023). The MillenniumTNG suite consists of numerous dark matter-only runs at various resolutions, two hydrodynamical runs named MTNG740 and MTNG185 in analogy with the IllustrisTNG naming scheme, and a collection of dark matter + massive neutrino runs over a range of resolutions and neutrino masses. In this article we consider only the hydrodynamical runs.
@@ 2.3 BAHAMAS
The BAHAMAS simulations use a $400\,\rm{kpc}$$\,h^{-1}$ box following cosmo-OWLS (Le Brun et al., 2014; McCarthy et al., 2014) and masses for the DM and baryonic particles a factor of $\sim 100$ larger than the highest resolution simulation considered here, IllustrisTNG300-1. As in Table 1, the gravitational softening length for the DM and stars in this simulation is identical (other than a small disagreement due to chosen cosmology) to that of IllustrisTNG300-3 at $\epsilon_{\rm{DM,\star}}=4\,\rm{kpc}$$\,h^{-1}$. This enables a comparison of offset measurements at a given softening length on similar mass scales, but with a different cosmology, feedback model and simulation code. For example the hydro method of BAHAMAS is based upon the smoothed particle hydrodynamics of gadget-3 whereas the hydro method utilized in IllustrisTNG is the moving voronoi mesh of AREPO. This provides a valuable opportunity to establish whether the findings of the offset measurements in this paper depend on some specific choice in the modelling which differs between TNG and BAHAMAS, or if despite the differences in simulation methodology the offset distributions are consistent.
@@ 3 Mass Maps
For each galaxy cluster of interest, we project the DM, stars and gas particles onto a grid with pixels of side length $0.25\,\rm{kpc}$, centered on the position of the most bound particle. We choose this pixel size to be smaller than the physical size of Hubble Space Telescope pixels for objects at redshifts $z\sim 0.5$, which is approximately $0.1^{\prime\prime}\sim 0.6\,\rm{kpc}$. The zoom box is chosen to have a side length of $50\,\rm{kpc}$ for the higher-resolution simulations (IllustrisTNG300-1, IllustrisTNG300-2 and MillenniumTNG), and a side length of $100\,\rm{kpc}$ otherwise. For the large scale measurements described in Sec. 4 (lensing and large scale gas) we use a $300\,\rm{kpc}$ zoom box at the same physical pixel size. Example mass maps for the highest-resolution simulation considered here, IllustrisTNG300-1, are shown in Fig. 1 along with the ellipses and various centroid identification methods described in the next section. The choice of pixel resolution and image field of view have the potential to impact offset measurements via their impact on centroid identification, and so we test the dependence of offset measurements on these parameters in Sec. 5.
In Fig. 1 and the remainder of this paper, we use the terminology “convergence radius”, a parameter which varies in definition (Power et al., 2003; van den Bosch & Ogiya, 2018) but also potentially from halo to halo in a simulation once a definition is fixed. Following Ludlow et al. (2019), the convergence radius may be estimated as $r_{\rm c}=0.055\,l$ where $l$ is the mean comoving interparticle separation of dark matter, $L_{\rm{box}}/N_{\rm{part}}^{1/3}$, with $L_{\rm{box}}$ the simulation box size and both parameters available in Table 1 for the simulations considered here. This considers the two-body relaxation between dark matter particles as the major mechanism affecting the convergence of halo central density profiles. Meanwhile, gravitational softening can also naturally “smooth” the matter distribution at a scale smaller than a few times $\epsilon_{\rm{DM,\star}}$. For simulations studied here, $\epsilon_{\rm{DM,\star}}$ values were specifically chosen to be of the same order as (and slightly smaller than) $r_{\rm c}$. Therefore, for simplicity in the rest part of this paper, we choose to take a conservative estimate of $r_{c}=2.8\epsilon_{\rm{DM,\star}}$, motivated by the “spline softening length" used by the GADGET code as the length scale above which pairwise forces become exactly Newtonian (Ludlow et al., 2019), but note that this choice is not impactful for the present analysis.
@@ 4 Centroid Identification
To measure offsets between the stellar mass distribution and a measure of the oscillation center, one must first establish a centroid for the stars and potentially other species like the DM or gas. In this section, we outline how we identify the center of the BCG in addition to various measures of the oscillation center. These are centroids obtained via the position of the most-bound particle in a cluster, source extraction via thresholding as a means of centroid fitting, and the centers of isophotes of cluster DM mass maps on scales relevant to cluster strong lensing.
One could also consider the usage of the cluster center of mass as a proxy for an oscillation center, however this probe is too sensitive to large-scale structure which is unrelated to the small-scale oscillation of the BCG in the core. In Ng et al. (2017), one method of determining the stellar centroid considered was the position of the “BCG”, defined therein by the peak of the i-band flux. Other choices considered included a weighted centroid of many member galaxies, the weighting being the i-band luminosity, and the peaks of heavily smoothed kernel density estimates (KDEs) of the galaxy luminosities. The authors observe that the BCG method exhibits by far the least scatter, therefore in the present work we do not consider the other methods, as when fitting centroids directly on small scales we are interested in the most faithful representation of an underlying “true” center.
@@ 4.2 Centroid Fitting
The offsets between centroids of different particle species, such as DM-stars or gas-stars offsets, can also be used as a proxy for the shape of the potential. To locate faithful centroids of these distributions, one can utilize the mass maps of Sec. 3. A centroid can be extracted from these maps by various means, for example by assuming a profile shape and fitting the profile using a minimization or posterior sampling scheme. One can also extract centroids without assuming a shape for the profile using an algorithm like shrinking spheres in 3D (eg. Schaller et al. 2015b), shrinking apertures in 2D, (eg. Ng et al. 2017), or by thresholding, whereby an underlying statistical variation is identified as a “background”, and any pixel a certain threshold above this background is classified as belonging to an object (eg Harvey et al. 2019). In the thresholding approach, the object pixels can then be divided into different sources, if present. A standard software package which performs this thresholding procedure is SourceExtractor (SE; Bertin & Arnouts 1996; Barbary 2016) and we choose to fit centroids of the DM, star and gas distributions using this method as opposed to profile fitting or convergent shrinking algorithms. This has the potential advantage of mimicking some systematics of the procedure applied to astronomical images to identify centroids of stellar and gas distributions, and we treat the DM in the same manner for consistency, noting that such a measurement would not be possible for real observations. This choice is consistent with the treatment of the stellar maps of Harvey et al. (2019), in which the stellar centroid was obtained via thresholding with SourceExtractor and the dark matter centroid/oscillation center was taken as the potential minimum (Sec. 4.1). Centers of stellar mass maps in this article are always identified using this thresholding procedure.

### v3

@@ Abstract
The distribution of offsets between the brightest cluster galaxies of galaxy clusters and the centroid of their dark matter distributions is a promising probe of the underlying dark matter physics. In particular, since this distribution is sensitive to the shape of the potential in galaxy cluster cores, it constitutes a test of dark matter self-interaction on the largest mass scales in the universe. We examine these offsets in three suites of modern cosmological simulations; IllustrisTNG, MillenniumTNG and BAHAMAS. For clusters above $10^{14}\,\rm{M_{\odot}}$, we examine the dependence of the offset distribution on gravitational softening length, the method used to identify centroids, redshift, mass, baryonic physics, and establish the stability of our results with respect to various nuisance parameter choices. We find that offsets are overwhelmingly measured to be smaller than the minimum converged length scale in each simulation, with a median offset of $\sim 1\,\rm{kpc}$ in the highest resolution simulation considered, TNG300-1, which uses a gravitational softening length of $1.48\,\rm{kpc}$. We also find that centroids identified via source extraction on smoothed dark matter and stellar particle data are consistent with the potential minimum, but that observationally relevant methods sensitive to cluster strong gravitational lensing scales, or those using gas as a tracer for the potential can overestimate offsets by factors of $\sim 10$ and $\sim 30$, respectively. This has the potential to reduce tensions with existing offset measurements which have served as evidence for a nonzero dark matter self-interaction cross section.
thanks: E-mail: roche@mit.edu
@@ 1 Introduction
The first evidence for the existence of a dominant and invisible matter component in the universe came from observations of the Coma galaxy cluster in 1933 (Zwicky, 2009), for which the kinematics of the galaxies within suggested an average density 400 times the inferred density from visible matter. The evidence for this dark matter (DM) has since accumulated to include for example the rotation curves of galaxies (Rubin & Ford, 1970; Ou et al., 2023), their escape velocities (Necib & Lin, 2022), the cosmic microwave background (CMB; Planck Collaboration et al. 2020), and gravitational lens modelling (Clowe et al., 2006; Sharon et al., 2020; Lin et al., 2022; Natarajan et al., 2024). The particle nature of DM remains one of the primary questions of modern astrophysics, but a cold, collisionless DM in tandem with a cosmological constant $\Lambda$ constitutes the current standard cosmological model (Weinberg et al., 2013; Vogelsberger et al., 2020). The $\Lambda$ cold dark matter (hereafter $\Lambda$CDM) cosmological paradigm has been remarkably successful in explaining the large scale structure of the universe via the Lyman-alpha forest (Evslin, 2017; Addison et al., 2018; Cuceu et al., 2019), the matter power spectrum and galaxy clustering (Reid et al., 2010; Eisenstein et al., 2011), the accelerated expansion of the universe (Riess et al., 1998; Perlmutter et al., 1999) and observations of the cosmic microwave background anisotropies and polarizations (Page et al., 2003).
@@
Galaxy clusters form at the densest nodes in the cosmic web, with the highest dark matter densities in their cores. These cores constitute ideal testing grounds to search for cored dark matter profiles, for example via strong lens modelling (Zitrin et al., 2012; Limousin et al., 2022), the internal kinematics of the brightest cluster galaxy (BCG; Kelson et al. 2002), weak lensing, and combinations of these probes (Newman et al., 2013). Evidence for both cored and cuspy cluster core profiles has been discovered using these methods (Biviano et al., 2023), with the source of the measured coring not yet clearly attributable to baryonic feedback or dark matter particle interactions alone. This uncertainty is due in part to the fact that baryonic effects have been demonstrated to both increase DM density slopes in the cores of clusters via adiabatic contraction (Schaller et al., 2015b), and decrease these slopes via dynamical friction (Inoue & Saitoh, 2011; Martizzi et al., 2012), galaxy formation processes (Laporte et al., 2012) and active galactic nucleus feedback (Ragone-Figueroa et al., 2012; Peirani et al., 2017).
@@
[…]
@@ 2.1 IllustrisTNG
The IllustrisTNG project (TNG; Springel et al. 2017; Pillepich et al. 2017a; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2021) is a suite of cosmological hydrodynamical simulations spanning various simulation box sizes, mass resolutions, and physics models based on the AREPO code (Springel, 2010). This project succeeds the Illustris project (Vogelsberger et al., 2014; Genel et al., 2014; Vogelsberger et al., 2014; Sijacki et al., 2015), improving upon the galaxy formation model (Vogelsberger et al., 2013) and expanding the scope to larger volumes and higher resolutions. The TNG simulation suite consists of TNG50, TNG100 and TNG300, characterized by their varying simulation box sizes of $35$, $75$ and $205\,\rm{cMpc}$$\,h^{-1}$ respectively.
@@ 2.2 MillenniumTNG
The MillenniumTNG project (MTNG; Hernández-Aguayo et al. 2023; Pakmor et al. 2023; Barrera et al. 2023) is the successor to the dark matter-only Millennium simulation (Springel et al., 2005), the first 10 billion particle simulation in a $500\,\rm{Mpc}$$\,h^{-1}$ box. MillenniumTNG combines the volume of Millennium with the hydrodynamical modelling [^2] With some small fixes in addition to changes relative to the TNG model, implemented to save memory in the much larger box (Pakmor et al., 2023). developed for the IllustrisTNG project (Weinberger et al., 2016; Pillepich et al., 2017b) and is run using an improved version of the AREPO code (Pakmor et al., 2015; Weinberger et al., 2020). The MillenniumTNG simulations involve high time resolution merger trees and direct light cone output, enabling detailed cosmological inference via comparisons to large scale surveys (Contreras et al., 2023; Kannan et al., 2023; Bose et al., 2023) and studies of the impact of baryons and massive neutrinos on weak lensing observations (Ferlito et al., 2023). The MillenniumTNG suite consists of numerous dark matter-only runs at various resolutions, two hydrodynamical runs named MTNG740 and MTNG185 in analogy with the IllustrisTNG naming scheme, and a collection of dark matter + massive neutrino runs over a range of resolutions and neutrino masses. In this article we consider only the hydrodynamical runs.
@@ 2.3 BAHAMAS
The BAHAMAS simulations use a $400\,\rm{cMpc}$$\,h^{-1}$ box following cosmo-OWLS (Le Brun et al., 2014; McCarthy et al., 2014) and masses for the DM and baryonic particles a factor of $\sim 100$ larger than the highest resolution simulation considered here, IllustrisTNG300-1. As in Table 1, the gravitational softening length for the DM and stars in this simulation is identical (other than a small disagreement due to chosen cosmology) to that of IllustrisTNG300-3 at $\epsilon_{\rm{DM,\star}}=4\,\rm{kpc}$$\,h^{-1}$. This enables a comparison of offset measurements at a given softening length on similar mass scales, but with a different cosmology, feedback model and simulation code. For example the hydro method of BAHAMAS is based upon the smoothed particle hydrodynamics of gadget-3 whereas the hydro method utilized in IllustrisTNG is the moving voronoi mesh of AREPO. This provides a valuable opportunity to establish whether the findings of the offset measurements in this paper depend on some specific choice in the modelling which differs between TNG and BAHAMAS, or if despite the differences in simulation methodology the offset distributions are consistent.
@@ 3 Mass Maps
For each galaxy cluster of interest, we project the DM, stars and gas particles onto a grid with pixels of side length $0.25\,\rm{kpc}$, centered on the position of the most bound particle. We choose this pixel size to be smaller than the physical size of Hubble Space Telescope pixels for objects at redshifts $z\sim 0.5$, which is approximately $0.1^{\prime\prime}\sim 0.6\,\rm{kpc}$. The zoom box is chosen to have a side length of $50\,\rm{kpc}$ for the higher-resolution simulations (IllustrisTNG300-1, IllustrisTNG300-2 and MillenniumTNG), and a side length of $100\,\rm{kpc}$ otherwise. The $50\,\rm{kpc}$ figure is chosen to be as small as possible while still containing sufficient shape information for reliable centering, since larger fields of view can produce larger offsets. See Sec. 5.5 for tests of the impact field of view and pixel size have on the offset distribution. For the lower resolution simulations, the $50\,\rm{kpc}$ mass maps are too noisy for reliable center identification, so we double the box width to gain constraining power. For the large scale measurements described in Sec. 4 (lensing and large scale gas) we use a $300\,\rm{kpc}$ zoom box at the same physical pixel size, chosen to have half-width approximately twice the Einstein radius of clusters at this mass scale (see Sec. 4.3). Example mass maps for the highest-resolution simulation considered here, IllustrisTNG300-1, are shown in Fig. 1 along with the ellipses and various centroid identification methods described in the next section.
In Fig. 1 and the remainder of this paper, we use the terminology “convergence radius”, a parameter which varies in definition (Power et al., 2003; van den Bosch & Ogiya, 2018) but also potentially from halo to halo in a simulation once a definition is fixed. Following Ludlow et al. (2019), the convergence radius may be estimated as $r_{\rm c}=0.055\,l$ where $l$ is the mean comoving interparticle separation of dark matter, $L_{\rm{box}}/N_{\rm{part}}^{1/3}$, with $L_{\rm{box}}$ the simulation box size and both parameters available in Table 1 for the simulations considered here. This considers the two-body relaxation between dark matter particles as the major mechanism affecting the convergence of halo central density profiles. Meanwhile, gravitational softening can also naturally “smooth” the matter distribution at a scale smaller than a few times $\epsilon_{\rm{DM,\star}}$. For simulations studied here, $\epsilon_{\rm{DM,\star}}$ values were specifically chosen to be of the same order as (and slightly smaller than) $r_{\rm c}$. Therefore, for simplicity in the rest part of this paper, we choose to take a conservative estimate of $r_{c}=2.8\epsilon_{\rm{DM,\star}}$, motivated by the “spline softening length” used by the GADGET code as the length scale above which pairwise forces become exactly Newtonian (Ludlow et al., 2019), but note that this choice is not impactful for the present analysis.
@@ 4 Centroid Identification
To measure offsets between the stellar mass distribution and a measure of the oscillation center, one must first establish a centroid for the stars and potentially other species like the DM or gas. In this section, we outline how we identify the position of the BCG in addition to various measures of the oscillation center. These are centroids obtained via the position of the most-bound particle in a cluster, source extraction via thresholding, and the centers of isophotes of cluster DM mass maps on scales relevant to cluster strong lensing. One could also consider the usage of the cluster center of mass as a proxy for an oscillation center, however this probe is too sensitive to large-scale structure which is unrelated to the small-scale oscillation of the BCG in the core.
@@ 4.2 Centroid Fitting
The offsets between centroids of different particle species, such as DM-stars or gas-stars offsets, can also be used as a proxy for the shape of the potential. To locate faithful centroids of these distributions, one can utilize the mass maps of Sec. 3. A centroid can be extracted from these maps by various means, for example by assuming a profile shape and fitting the profile using a minimization or posterior sampling scheme. One can also extract centroids without assuming a shape for the profile using an algorithm like shrinking spheres in 3D (eg. Schaller et al. 2015a), shrinking apertures in 2D, (eg. Ng et al. 2017), or by thresholding, whereby an underlying statistical variation is identified as a “background”, and any pixel a certain threshold above this background is classified as belonging to an object (eg Harvey et al. 2019). In the thresholding approach, the object pixels can then be divided into different sources, if present. A standard software package which performs this thresholding procedure is SourceExtractor (SE; Bertin & Arnouts 1996; Barbary 2016) and we choose to fit centroids of the DM, star and gas distributions using this method as opposed to profile fitting or convergent shrinking algorithms. This has the potential advantage of mimicking some systematics of the procedure applied to astronomical images to identify centroids of stellar and gas distributions, and we treat the DM in the same manner for consistency, noting that such a measurement would not be possible for real observations. This choice is consistent with the treatment of the stellar maps of Harvey et al. (2019), in which the stellar centroid was obtained via thresholding with SourceExtractor and the dark matter centroid/oscillation center was taken as the potential minimum (Sec. 4.1). Centers of stellar mass maps in this article are always identified using this thresholding procedure.
@@
We refer to the center of the highest-flux source in the stellar mass map as the position of the “BCG”, however it has been demonstrated (van den Bosch et al., 2005; Skibba et al., 2011; Lange et al., 2018) that $\sim 30\%$ of clusters have central galaxies which are not their brightest member galaxy. Therefore the offsets in this paper are most accurately described as offsets between some measure of the oscillation center and the brightest source within $\lesssim 25\,\rm{kpc}$ of the potential minimum (set by half the side length of the zoom box, and therefore this number is variable). This source may be the BCG or a central galaxy which is not the brightest member, but for simplicity we maintain the terminology of a “BCG offset”. Note that various other stellar centroid measures exist, for example in Ng et al. (2017), stellar centroids such as the peak of the i-band flux, an i-band luminosity-weighted centroid of many member galaxies, and the peaks of heavily smoothed kernel density estimates (KDEs) of the galaxy luminosities were each considered, with the peak of the $i$-band flux exhibiting the least scatter.